### v1

# Imperceptible Rhythm Backdoor Attacks: Exploring Rhythm Transformation for Embedding Undetectable Vulnerabilities on Speech Recognition

Abstract
Speech recognition is an essential start ring of human-computer interaction, and recently, deep learning models have achieved excellent success in this task. However, when the model training and private data provider are always separated, some security threats that make deep neural networks (DNNs) abnormal deserve to be researched. In recent years, the typical backdoor attacks have been researched in speech recognition systems. The existing backdoor methods are based on data poisoning. The attacker adds some incorporated changes to benign speech spectrograms or changes the speech components, such as pitch and timbre. As a result, the poisoned data can be detected by human hearing or automatic deep algorithms. To improve the stealthiness of data poisoning, we propose a non-neural and fast algorithm called Random Spectrogram Rhythm Transformation (RSRT) in this paper. The algorithm combines four steps to generate stealthy poisoned utterances. From the perspective of rhythm component transformation, our proposed trigger stretches or squeezes the mel spectrograms and recovers them back to signals. The operation keeps timbre and content unchanged for good stealthiness. Our experiments are conducted on two kinds of speech recognition tasks, including testing the stealthiness of poisoned samples by speaker verification and automatic speech recognition. The results show that our method has excellent effectiveness and stealthiness. The rhythm trigger needs a low poisoning rate and gets a very high attack success rate.
keywords:
Backdoor attacks , Speech recognition , Rhythm transformation , Neural vocoder
journal: Neurocomputing
[…]
1 Introduction
Speech recognition systems are critical components of human-computer interaction, [1] which enables machines to recognize human identity or vocal commands. Speech recognition models are usually trained by machining learning methods and need abundant supervised utterance datasets and precious computational resources. Under special circumstances, some companies entrust their sensitive speech recognition datasets to third-party training platforms to reduce training expenses.
However, recent research found that exposing classification datasets to malicious training developers may make the deep neural network (DNN) models vulnerable [2]. In some training procedures ($e.g.$, data collection and preparation and model training), the attackers can manipulate the behavior of speech recognition systems by embedding backdoors to DNN models and cause an extreme security risk. The backdoor adversaries poisoned the model to learn the benign and attacker-specific tasks by implanting the backdoor into the target. The adversaries usually generate poisoned samples and alter their ground truth labels with designed triggers for the poisoned task. For inputs containing no trigger, the victim model behaves normally as its clean parallel model. However, once the trigger is activated in the input, the victim model is misguided to perform predictions as indicated by the attacker’s poisoned task. It is difficult to distinguish the backdoored model from its clean version by simply checking the test accuracy with the test dataset.
Most of the backdoor attack methods are developed in computer vision tasks and text classification at present [3, 4, 5, 6, 7]. These methods usually treat noisy pixel patterns and extra phrases as triggers. Motivated by these, the study of backdoor attacks in speech recognition imitates these methods, whose triggers are ultrasonic sound, hidden noisy shrill, monotone sound, and some time-frequency mask of the spectrogram [8, 9, 10, 11, 12, 13, 14]. In latest research, the trigger in speech starts shifting to the components of the speech, such as pitch boosting and timbre conversion [15, 16, 17, 18]. However, the extra noisy clips destroy speech quality and make the trigger unconcealed. Besides, the pitch and timbre triggers have the potential to be automatically detected. According to voice disentanglement research [19, 20], four main speech components are considered important: rhythm, content, timbre, and pitch. However, the transformation of pitch and timbre can be detected by Automatic Speech Recognition (ASR) and speaker verification systems (SVS); the pitch boosting can also be detected by the YIN algorithm [21].
Can the backdoor trigger in speech recognition avoid automatic detection and sustain naturalness and speech quality? In this paper, we give a positive answer, and we target rhythm as a trigger from the perspective of the components. The rhythm is mainly related to the speed of each syllable [22]. We propose a non-neural and fast algorithm called Random Spectrogram Rhythm Transformation (RSRT) to generate poisoned samples whose rhythms are transformed. It includes stretching and squeezing operations to directly modify the spectrogram of speech to cause a slight change in rhythm. The poisoned spectrogram is then reconstructed into speech using a neural network vocoder, ensuring the converted speech’s naturalness and intelligibility. Numerous studies have shown that the neural network vocoder exhibits good generalization performance [23, 24] for various modified spectrograms. We mainly focus on the keyword spotting (KWS) task and text-independent speech emotion recognition (TSER) task in our work because the slight change of rhythm does not destroy the content and emotion. Finally, we conducted two evaluation metrics on poisoned samples to verify the consistency of speech components. The experiment results demonstrate that the rhythm trigger gains a high attack success rate with a very low poisoning rate. Our contributions can be summarized as follows:
  • 1.
    We designed a non-neural rhythm transformation poisoning pipeline containing RSRT. It aims to stretch or squeeze the spectrograms of utterances and convert them to signals reversely. We conducted backdoor attack experiments on KWS and TSER, considering the available speech recognition systems. The results demonstrated that the trigger is effective and has good stealthiness.
  • 2.
    We conducted three kinds of evaluation experiments to prove the good stealthiness of our proposed trigger. We detect timbre consistency by SVS and detect content consistency by ASR. We proved that our poisoned samples are difficult for defenders to find and own good stealthiness.
The rest of this paper is structured as follows. In Section 2, we briefly introduce backdoors and speech recognition. In Section 3, we illustrate backdoor method motivation and attacking theory. we elaborately describe the main stages of RSRT. In Section 4, we show the results of attack effectiveness and stealthiness evaluation. Finally, we conclude this paper in Section 5 at the end.
2 Related Works
2.1 Speech Recognition
2.2 Backdoor Attacks
Backdoor attacks [25, 26] are a series of poisoning methods that aim to bury a backdoor to the model parameters. The attackers usually design specific target labels and guarantee the model will learn the relationship between labels and triggers. Depending on whether training data is poisoned, backdoor attacks can be categorized into poisoning-based attacks and clean-based attacks. Although the poisoning-based attacks enhance the effectiveness of backdoor attacks, the poisoned samples are vulnerable to avoiding detection by defenders [27, 28, 29]. To avoid notice of defender while remaining attacks, Turner et al. [30, 31] firstly explored clean-label backdoor attacks that aim to control the model’s behavior by an adversary without poisoning input samples during inference. From the perspective of attackers, we will explain various attack methods.
2.2.1 Poisoned-Label and Clean-Label Attacks
Poisoned and clean-label attacks usually happen in classification models that learn weight optimization by many pairs of samples and labels. The poisoned-label attackers modify both samples and labels, while the clean-label attackers only modify the samples. Chen et al. blended an attacker-specific Hello Kitty pattern as a trigger with images [32]. Eykholt et al. constructed distributed and spread images as triggers to mask sign images [33]. Sarkar et al. [34] utilized existent facial characteristics as triggers. All these modifications are connected to target labels. Turner et al.[35] first explored a method that leveraged adversarial perturbations or generative models to modify some benign images for clean-label attacks.
2.2.2 Sample-Agnostic and Sample-Specific Attacks
The poisoned images are commonly generated by attacker-specified generators with specific trigger patterns. Sample-agnostic attacks are widespread backdoor attacks in which all poisoned samples share the same trigger. However, this characteristic can be easily exploited for backdoor defense [27, 36]. Nguyen et al. [37] proposed applying visible triggers to be contained in different poisoned samples, which is the first sample-specific backdoor attack method. This method embedded trigger patterns in the edge structure of poisoned images, broke many defense assumptions, and led to good effectiveness on many backdoor defenses. Therefore, sample-specific attacks cause a serious security threat and are worth deeper investigations.
2.2.3 Semantic and Non-Semantic Attacks
The attackers may utilize the semantic information in the images and relate semantic regions with target labels for backdoor attacks, and this method is a semantic attack. Bagdasaryan et al. [38, 39] first explored that certain features are $e.g$, green cars or white birds can be related to attacker-specified target labels. It is noted that the semantic information is ”bird” or ”car.” Accordingly, the victim model will wrongly classify images containing pre-defined semantic information without any digital modification during inference.
2.2.4 All-to-One and All-to-All Attacks
According to the type of target labels, poisoning-label backdoor attacks can be categorized as all-to-one attacks and all-to-all attacks. All-to-one attacks assume that poisoned samples share the same target label, regardless of their original labels. On the contrary, poisoned samples may possess different target labels within all-to-all attacks. For example, Doan et al. [40, 41] applied label shifting function assigning as $LS(y)=(y+1)\mod T$. It noted that $T$ is a controllable integer by attackers, and $y$ is label.
2.2.5 Digital and Physical Attacks
Previous attacks usually occur entirely in digital space, and attackers implement triggers by coding, which belongs to totally digital attacks. Physical attacks are involved in generating poisoned samples in physical space. Chen et al. [32] firstly applied a pair of glasses as a physical trigger for poisoning a face recognition system originating in a camera. Wenger et al. [42, 3] also found that colored dot stickers can be treated as physical triggers. However, Li et al. [43] demonstrated that digital triggers are sensitive to space transformation, such as rotation and shrinkage, which will change the location and form of triggers in poisoned samples. Based on this knowledge, they proposed a transformation-based attack enhancement that keeps the improved attacks effective in the physical world.
3 Methods
3.1 Motivation
According to the backdoor attack principle, we wish that poisoned speech samples are stealthy while facing automatic or human-hearing detection. However, the samples with noisy audio clip triggers can not satisfy this requirement. Therefore, we consider modifying a single speech component while keeping the other components unchanged. The components are shown in Figure 1.
Figure 1: Backdoor attacks by changing speech components.
In [17, 16], the stealthiness evaluation has demonstrated that tiny modifications in timbre and pitch do not influence speech naturalness and intelligibility. However, deep speech systems such as SVS can find the modification.
In this paper, we aim to treat the rhythm as the backdoor trigger. The rhythm is highly correlated with the duration of each syllable. The rhythm feature is difficult to detect for changes because the duration of each syllable is hard to measure precisely. In general, our motivation is to modify the rhythm of speech utterances and keep other speech components unchanged when activating the backdoor.
Table 1: The definition of backdoor description symbols.
Notation Description
$f_{\theta}$ speech classifier learned from benign dataset
$f_{\theta^{\prime}}$ speech classifier learned from poisoned samples
$\mathcal{X}\times\mathcal{Y}$ domain space of inputs and labels
$\mathcal{F}_{t}:\mathcal{X}\rightarrow\mathcal{X^{*}}$ backdoor input trigger
$\mathcal{F}_{y}:\mathcal{Y}\rightarrow\mathcal{Y^{*}}$ label shifting function
$D,D_{e}$ benign training and test dataset
$D_{r}$ selected subset from benign dataset
$D_{s}$ poisoned subset from selected subset
$D_{p}$ poisoned dataset that contains poisoned and benign samples
$L_{(x,y)}$ training objective that is training on dataset $\{(x,y)\}$
3.2 Preliminaries
3.2.1 Neural Vocoder
Neural vocoder is a neural network that converts spectrograms to speech signals and exhibits excellent generalization performance, which encourages the use of it for reconstructing stretched and squeezed spectrograms. The vocoder used in our experiment is HiFi-GAN [44], which includes a generator and discriminator. The pre-trained generator is applied for conversion.
Figure 2: The proposed attack pipeline via RSRT. The pipeline consists of three main stages, including (a) the Attack Stage, (b) the Training Stage, and (c) the inference Stage. The RSRT is the core of the attack stage, containing four steps (i.e., VAD, RSRT, vocoder conversion, and silence concatenation). 1) We use VAD to extract and locate active speech regions for effective attacks. 2) We choose a set of rhythm transformation hyper-parameters and use RSRT to stretch or squeeze utterances, making the rhythm migration. 3) The rhythm-migrated speech is converted back to signal with a pre-trained neural vocoder, maintaining consistency in speech content and timbre. 4) To make the poisoned speech similar to normal speech, we concatenated silence at the beginning and end, ensuring that the duration of the poisoned speech matches the original speech.
3.2.2 Threat Model
: This paper concentrates on poisoning-based backdoor attacks. And there are some basic principles in this scenario. The adversaries can only modify the open-access training dataset to a poisoned dataset. The victim models will be trained on the poisoned dataset, and the user will deploy the models to the working environment. Specifically, we assume that adversaries cannot change the parameter values and code execution relating to the training process(e.g., loss function, learning schedule, or the resulting model.
3.2.3 The Goal of Adversary
: The attacker’s goals primarily include stealthiness, effectiveness, and robustness. Stealthiness requires that backdoor attacks have the ability to escape human examination and machine detection. Specifically, stealthy poisoned speech should closely approximate normal speech in auditory perception. Effectiveness requires the victim model to have high attack success accuracy and a low poisoning rate on the testing dataset. It is noted that although some methods achieve very high attack success rates, they often require a concerning proportion of poisoned samples. This configuration may lead to bad stealthiness. Robustness requires that backdoor attacks behave well on simple detection means and remain effective under more difficult settings, such as adaptive defenses and physical-world settings.
3.2.4 Poisoning-Based Backdoor Attacks Pipeline
We first illustrate backdoor attacks by the notions and their definition of backdoor in Table 1. We denote the classifier $f_{\theta}:\mathcal{X}\rightarrow\mathcal{Y}$, where $\theta$ signifies model parameters, $\mathcal{X}\in R^{T,C}$ being the instance space, and $\mathcal{Y}=[1,2,...,K]$ being the label space. The $T,C$ represent the sequence length and channel number. Let $\mathcal{F}_{t}:\mathcal{X}\rightarrow\mathcal{X}$ indicates the attacker-specified trigger function and $\mathcal{F}_{y}:\mathcal{Y}\rightarrow\mathcal{Y}$ indicates label shifting function. Before attacking, the clean training dataset is prepared that is signified as $D=\{(x_{i},y_{i})\}^{N}_{i=1}$, then, the attacker will design the poisoned subset that is conducted by $D_{s}=\{\mathcal{F}_{t}(x_{j}),\mathcal{F}_{y}(y_{j})\}^{M}_{j=1}$ where the replaced subset is $D_{r}=\{(x_{k},y_{k})\}^{M}_{k=1}$. Finally, the poisoned dataset is mixed by $D_{p}=(D-D_{r})\cup D_{s}$. Backdoor attacks request the model to optimize $f_{\theta}$ by following the training objective.
$L_{(x,y)\in D_{p}}=\mathop{\arg\max}\limits_{\theta}p(y|f_{\theta}(x))$ (1)
This objective leads the model to correctly classify the benign samples $x\in\mathcal{X}$ to their ground true labels and the poisoned samples $x^{*}=\mathcal{F}_{t}(x)\in\mathcal{X^{*}}$ to target labels respectively. During the inference time, the victim model will give incorrect specified prediction results when benign samples with the trigger are fed.
Figure 3: The illustration of rhythm transformation. (a) denotes the process of stretching algorithm. Some frames are copied in the next places of the original index, while the other frames are retained. (b) denotes the process of squeezing the algorithm. A part of the frames and their next frames are selected to form new frames by double linear weight sum. (c) and (d) respectively show the speech spectrograms squeezed to $1/2$ times and $2/3$ times and stretched to 1.3 times to 2.0 times.
3.3 Attack via Random Spectrogram Rhythm Transformation
The typical speech triggers borrow methods straightforwardly from image backdoor attacks. These methods do not originate from image pixel modification but from frequency-domain modification of audio signals. For example, PIBA trigger [9] adds a short noise clip to the signal. JingleBack trigger [45] applies pitch shift and distortion to the signal. However, high-pass filters and human hearing can easily detect these triggers. The triggers of modifying speech components retain the naturalness of poisoned utterances without adding any external sounds or making complex distortion. PBSM[17] and VSVC[16] proposed to change the pitch and timbre. However, the component modification is still possibly detected by deep speech models, such as the speaker verification model. To tackle this problem. We turn our attention to another component, rhythm. The rhythm refers to the speed of each speech syllable, representing the pace of the spoken language. We propose to change each syllable speed by a simple spectrogram frame-level algorithm. Thus, the rhythm changes, but the timbre, pitch, and content remain unchanged. The algorithm is used in the attack stage via random spectrogram rhythm transformation (RSRT).
The poisoning-base attack pipeline via RSRT focuses on three stages, as shown in Figure 2, including (a) the Attack stage, (b) the Training stage, and (c) the Inference stage. We mainly describe the theory of the attack stage as shown in Figure 2(a). It includes (1)Voice Active Detection (VAD), (2)RSRT, (3)Vocoder conversion, (4)Silence concatenation. The RSRT algorithm is the key operation. In the attack stage, it firstly extracts active speech regions by energy-based VAD [46], then performs frame-level stretching or squeezing to the active spectrogram regions. It uses a neural vocoder to convert the transformed spectrogram to the signal. Finally, it reassembles them with silent clips to form the same speech length as the original. The final operation helps the poison utterances behave like benign utterances to avoid simple machine detection defense. Then, we will describe each step in detail.
3.3.1 Voice Active Detection
We used energy-based Voice Active Detection (VAD) to discriminate between silent regions and active voice regions. Given a spectrogram $X^{D,T}=\{x_{i}\in R^{D}|i=1,2,...,T\}$, the average energy of every frame will be calculated as follows,
$E_{x}=\{\frac{1}{D}\sum\limits^{D}_{j=1}x_{i}|i=1,2,...,T\}$ (2)
We set a threshold $S_{e}=\mu*\max(E_{x})$ equal to $\mu$ times the maximum value of $E_{x}$. It is noted that $\mu<1$. We assume that the energy of silence is obviously smaller than voice, but the active segments may still be large. Thus, the coefficient is set to close to one time. To avoid detecting some short recording noise and shoddy sound to speech, we decide the vocal continuous frames whose average energy values are upper than the threshold as voice segment $X_{voi}$ as follows,
$X_{voi}=\{x_{i}|E_{x}(i)\geq S_{e},i=m+1,m+2,...n\}$ (3)
The $X_{voi}$ represents an active region in an utterance. We show an example of the VAD result in Figure 4.
Figure 4: The VAD result. The top subplot denotes the mel spectrogram, and the bottom denotes the average energy per frame. The red box highlights the non-voice portion.
The blue curve in the subplot below Figure 4 indicates the value of the average energy $E_{x}$. We have marked the position of the speech $X_{voi}$ with red coordinates. The segment highlighted by the red box shows a different property to pure speech, observed as non-voice things, such as noise and recording disorder.
3.3.2 RSRT Methods
The RSRT algorithm aims to change speech rhythm and connect different rhythms with target labels, including stretching and squeezing operations. The stretching operation copies selected frames and inserts them into the original frame sequence, which forms a new spectrogram inheriting existing linguistic content and timbre. The stretching operation is shown in Figure 3(a). The squeezing operation uses a bilinear downsampling algorithm [47] to generate news frames from single syllables. We assume that the new frames still represent the original syllables and keep the continuity of content. The squeezing operation is shown in Figure 3(b). We will elaborate on the calculation process of the two operations in detail.
(1) Stretching: Given a spectrogram $X^{D,T}=\{x_{i}|i=1,2,...,T\}$ composed of frames, where D is the number of frequency bins, and T is the number of frames $x_{i}$. It assumes that the stretched spectrogram is $Y_{s}=\{y_{j}|j=1,2,...,T_{s}\}$. Given a scale ratio parameter $\gamma_{s}$ and frame repetition count $\sigma_{s}$, each new frame can be derived by following algorithm 1:
Algorithm 1 Stretching algorithm
$\gamma_{s}\in[0,1]$, $\sigma_{s}>0$, $X=\{x_{i}|i=1,2,..,T\}$.
2:$T_{s}\leftarrow\lfloor(T+\gamma_{s}*T*\sigma_{s})\rfloor$
3:$Y_{s}\leftarrow\{y_{j}=0|j=1,2,...,T_{s}\}$.
4:$Old\_index\leftarrow\{i|i=1,2,...,T\}$
5:$m\leftarrow 1$
6:while $!Empty(Old\_index)$ do
7:     $d\leftarrow RandomChoice(Old\_index)$
8:     $r\_ind\leftarrow\{d,d+1,d+2,...d+\sigma_{s}\}$
9:     $Y_{s}\left[r\_ind\right]\leftarrow X[d]$
10:     $m\leftarrow m+1$
11:     if $m\geq T_{s}$  then
12:         $break$
13:     end if
14:end while
15:return $Y_{s}$
In the calculation, the algorithm selects the indexes of a $\gamma_{s}$ proportion of frames, which are repeated by $\sigma_{s}$ times. Thus, the length of frames in $Y_{s}$ is extended to $\lfloor(T+\gamma_{s}*T*\sigma_{s})\rfloor$.
(2) Squeezing: It is worth mentioning that maintaining the continuity of the spectrogram is crucial in our proposed squeezing algorithm. As seen in the stretching algorithm, each frame of the $Y_{s}$ spectrogram maintains continuity. Therefore, when choosing those frames to squeeze, we must maintain the same continuity by avoiding randomly choosing the indexes but following an arithmetic sequence that starts from 0, ends at $T$, and has a common difference of $\phi_{c}$. We believe that using a bilinear downsampling algorithm can maintain the continuity of the speech spectrogram.
Given a spectrogram $X$ like above, it is assumed that the squeezed spectrogram is $Y_{c}=\{y_{k}|k=1,2,...,T_{c}\}$. Given a common difference parameter $\phi_{c}$ and bilinear weight $w$, each new frame can be derived by following algorithm 2:
Algorithm 2 squeezing algorithm
1:$\phi_{c}\geq 2$, $X=\{x_{i}|i=1,2,..,T\}$.
2:$down\_index\leftarrow\{D_{k}=\phi_{c}*(k-1)+1|k=1,2,...,N.D_{k}\leq T\}$triangleright arithmetic sequence of indexes
3:$down\_next\leftarrow down\_index+1$
4:$X\left[down\_index\right]\leftarrow(1-w)*X\left[down\_index\right]+w*X\left[down\_next\right]$
5:$nd\leftarrow\left[\{1,2,...,T\}-down\_next\right]$
6:$X\leftarrow X\left[nd\right]$
7:return $X$
In the spectrogram squeezing algorithm, we selected a series of indexes of some frames in the order of an arithmetic sequence, whose common difference is $\phi_{c}$. Then we replaced these frames at the positions of these indexes with the sum of their bilinear weights with the frame following them. The length of new $X$ is about scaled to $(1-\frac{1}{\phi_{c}})*T$. The examples of RSRT are shown in Figure 3. We set a series of parameters to stretch a clean utterance from 1.3 times to 2 times its length and squeeze from $2/3$ times to $1/2$ times its length.
3.3.3 Vocoder Conversion
The conversion stage aims to convert the stretched or squeezed spectrogram to a signal, maintaining consistency in speech content and timbre. It’s worth noting that the signal parameters(e.g. sample rate, hop length) for spectrogram extraction used in both the vocoder training and speech recognition training need to be equal, ensuring that the speech generated by our trigger can be restored without lack of sample rate and quality.
3.3.4 Silence Concatenation
It is observed that many utterances in the dataset have non-voice things like noise or recording disorder, as Figure 4 shows. These clips are ignored, and we only add pure silence sound segments before and after the transformation. In addition, we kept the duration of the poisoned samples the same as the original samples.
4 Experiments and Results
We evaluate the proposed attack pipeline on KWS and TSER experiments. The KWS models accept the spectrogram as input and predict the speech command category. The TSER models accept the same input and output the speech emotion category.
4.1 Experimental Setting for KWS Task
Dataset: We use the Google Speech Commands Dataset v2 [48] as our experimental materials. The dataset contains more than 65,000 audios of 35 categories of keywords that can be used for keyword-spotting devices. In our experiments, 23,682 audios of 10 categories(“yes,” “no,” “up,” “down,” “left,” “right,” “on,” “off,” “stop,” and “go”) are used. We divide the dataset into the training, validation, and test sets in a ratio of 95:5:5, where the validation set belongs to the training set. The poisoned samples only exist in the training set.
Victim models: Our experiments were performed on the following four KWS networks: Resnet-34 [49], Attention-LSTM [50], KWS-ViT [51], EAT-S [52], they behave excellent classification performance on the keyword spotting task.
Baseline and Attack Setup: We compare our attack with the latest speech backdoor attacks. They are listed as follows. (1) backdoor attack with pixel pattern(BadNets) [3], (2) position-independent backdoor attack(PIBA) [9], (3) dual adaptive backdoor attack(DABA) [13], (4) ultrasonic voice as trigger(Ultrasonic) [10], (5) pitch boosting and sound masking(PBSM) [17], and (6) voiceprint selection and voice conversion(VSVC) [16].
In our RSRT method, We set $\sigma_{s}=1$ and $\gamma_{s}=\{0.5,1.0\}$ for stretching the original duration to $\{1.5,2.0\}$ times. We set $\phi_{c}=\{2,3\}$ and bilinear weight $w=0.6$ for squeezing the original duration to $\{\frac{1}{2},\frac{2}{3}\}$ times. In the stage of VAD, the threshold coefficient is set to 0.85.
Training Setup: We trained all the victim models with the same hyper-parameters. The batch size is 64. The weights are optimized by Adam optimizer with a learning rate of 1e-4 and cross-entropy loss function. We trained 30 epochs to make all models converge.
Table 2: Evaluation of backdoor attack methods on KWS task. Each item shows the $AV/ASR/PN$ in the table.
KWS Models
Trigger Resnet-34 Attention-LSTM KWS-ViT EAT-S
BadNets 1.97/96.48/300 2.04/97.05/300 2.15/96.66/350 2.68/96.67/350
PIBA 2.68/94.21/300 2.92/93.58/350 3.15/94.62/350 3.61/93.59/350
DABA 3.65/93.25/450 4.21/92.52/400 3.91/92.55/450 4.55/93.45/450
Ultrasonic 1.24/95.42/400 1.56/96.41/400 1.72/93.57/450 1.64/95.64/450
PBSM 0.78/99.95/300 0.82/99.85/300 0.97/99.76/400 0.69/99.85/400
VSVC 0.51/99.98/250 0.50/99.97/250 0.67/99.92/300 0.56/99.93/250
RSRT(Stretch) 0.48/99.97/150 0.60/99.97/150 0.65/99.94/200 0.47/99.95/200
RSRT(Squeeze) 0.61/99.93/150 0.55/99.93/200 0.51/99.91/150 0.61/99.96/200
Table 3: Evaluation of backdoor attack methods on TSER task. Each item shows the $AV/ASR/PN$ in the table.
TSER Models
Trigger AST SER-AC SER-CNN
BadNets 3.78/92.14/550 4.20/93.15/500 3.82/94.15/500
PIBA 4.05/95.62/500 4.65/96.14/500 4.17/97.15/500
DABA 3.64/98.65/450 4.02/98.72/400 4.12/98.56/400
Ultrasonic 2.67/97.82/350 2.92/97.68/400 3.01/96.92/400
PBSM 0.97/99.58/450 0.96/99.67/400 0.98/99.72/400
VSVC 0.98/99.94/350 0.92/99.97/400 0.93/99.94/400
RSRT(Stretch) 1.09/99.87/250 1.31/99.89/250 1.61/99.25/200
RSRT(Squeeze) 1.22/99.86/200 1.56/99.69/200 1.42/99.25/250
4.2 Experimental Setting for TSER Task
Dataset: We used two speech emotion datasets, Emotional Speech Dataset (ESD) [53] for the TSER task. The ESD dataset consists of 350 parallel utterances spoken by 10 native English and 10 native Chinese speakers and covers 5 emotion categories (neutral, happy, angry, sad, and surprise). We only use ESD samples from a single language for training.
Victim models: Our experiments were performed on signal process deep neural models. We choose AST [54], SER-AC [55] and SER-CNN [56]. These models use only signal information to learn emotional classification and reach effective performance.
Baseline and Attack Setup: We set the same proposed attacking configuration as the KWS task.
Training Setup: We trained all the victim models with AST’s hyperparameters. The batch size is 12. The weights are optimized by Adam optimizer with a learning rate of 1e-5 and cross-entropy loss function. The learning rate is cut into half every epoch after the 2nd epoch. We trained 30 epochs to make all models converge.
4.3 Evaluation Metrics
Evaluation metrics reflect the stealthiness and effectiveness of the proposed trigger.
Attack Metrics: [57] We mainly consider attack success rate (ASR), accuracy variance (AV), and poisoned number (PN). The backdoor attackers aim to achieve the highest possible attack success rate with minimal poisoning rate and accuracy variance.
Stealthiness Metrics: To ensure the speech components except rhythm are consistent. We use timbre consistency rate (TCR) and word error rate(WER) [58] to measure the ratio that poisoned samples keep the timbre and content consistent. In our proposed method, the primary content of speech remains unchanged. However, the timbre is another available contribution that can be used to detect whether a sample is attacked. Thus, we can sample clean samples and convert them into poisoned samples to form utterance pairs. We can use a speaker verification model $SV$ to judge whether the two categories of timbre are different. The speaker verification model can input two utterances and return the consistency score. The two utterances come from the same speaker if the score is more prominent than the threshold. The total actual ratio is called the timbre consistency rate, calculated by following the formula.
$e_{1}=SV(x),e_{2}=SV(y)$ (4)
$sc=E(e_{1}*e_{2}^{T})$ (5)
$Score(x,y)=\begin{cases}1&sc>=threshold\\ 0&sc<threshold\end{cases}$ (6)
$TCR=\frac{\mathcal{I}(Score(x_{i},\mathcal{F}_{t}(x_{i})))}{N_{c}}\in D_{e}$ (7)
On the other hand, WER is a common metric to compare the difference in speech content between predicted and ground truth words. We use the latest SV system ERes2Net [59] for TCR evaluation, and the accepted score threshold is set to $0.70$. We also used paraformer [60] for WER evaluation with its open-source code.
4.4 Main Results.
4.4.1 Attack Results
Tables 2 and 3 show the main results of backdoor attack evaluation on KWS and TSER tasks. Table 4 shows the timbre consistency evaluation result. We randomly select 500 clean samples in the test dataset and generate poisoned samples with all triggers. We compare the main evaluation metrics, which include AV, ASR, and PN. In the experiments, we found that all the ASR values will increase with the increase of PN values, ultimately approaching $100\%$. Thus, the tables show all the highest ASR values in the tables with the highest PN. The stretching ratio is $2.0$, and the squeezing ratio is $0.5$ in Tables 2 and 3. Then, we will first analyze our method.
In the proposed stretching and squeezing results, the PN values are less or equal to 200 in all cases of the proposed method. However, the PN values of other methods are all greater than 200. It is noted that 500 poisoned samples equals about $2.9\%$ and $0.4\%$ poisoning rate. In conclusion, our proposed method has excellent stealthiness due to the low poisoning rate and high attack success rate. Because our proposed method changes the speech rhythm while keeping the content, timbre, and emotion unchanged. The AV results are also smaller than most of the baselines.
Next, we analyze the evaluation results of the baseline methods. The triggers can be classified in two ways. One is adding noisy clips to clean speech utterances or incorporating spectrograms using special patterns. These triggers are BadNets, PIBA, DABA, and Ultrasonic as shown in Figure 5(d-e). On the other hand, triggers (PBSM and VSVC), as shown in Figure 5(e-f), change a single speech component and keep excepted components unchanged. The BadNets trigger has a very high AV and a decent ASR in two tasks. That is to say, it has good effectiveness but decent stealthiness. Because the trigger adds a tiny pixel-level single pattern to a benign spectrogram, it can not influence the whole recognition of the utterance.
PIBA, DABA, and Ultrasonic triggers make complex incorporation into spectrograms. Due to extreme spectrogram modification in specific time-frequency domain. The AV and ASR values are both high. However, these methods require a high poisoning number for the model not to recognize them as noise. During the early stages of neural network training, they may still be regarded as noise, leading to disruption of emotion and speech quality, thus affecting the learning ability of the model from the outset and ultimately resulting in poorer training outcomes.
The VSVC trigger can change the timbre of speech to an attacker-specified target timbre in the training set while the content and rhythm stay the same. With non-parallel and GAN training, the voice-converted poisoned utterances are of good quality, so the AV values are both great. This trigger associates speech of specific timbre with target labels, while the KWS and TSER tasks are independent of timbre recognition. Therefore, this method also has high ASR results. However, training a voice conversion neural network is time-consuming, and not robust enough. Because the defender can first back up some utterances in their database as ”timbre private keys.” Before the training phase, these can be compared with all data in the dataset(equivalent to public keys) with a speaker verification system to detect the presence of voice conversion backdoor attacks.
The PBSM trigger boosts the pitch of speech and masks the boosted voice with a masking sound to form a peak sound as a poisoned sample. This operation also slightly degrades the speech’s quality, which leads to low AV values and a good poisoning rate. However, the change of pitch is easy to be detected by the signal algorithm. Thus, Our proposed method generally exhibits optimal performance in both stealthiness and effectiveness.
4.4.2 Stealthiness Evaluation Results
Table 4 shows all the evaluations of speech backdoor methods’ TCR and WER. We also show mel spectrograms with different triggers in Figure 5. As shown in Table 4. We use baseline and proposed triggers in the evaluation to form an utterance pair. Then, we test whether the clean and poisoned ones can be derived from the same speaker and content by a speaker verification system and automatic speech recognition system. Our proposed method can retain the content and timbre well, while the TCR and WER are close to $1$ and 00 in stretching or squeezing operations. The VSVC trigger gets very low TCR and high TCR because the timbre converts completely. The PBSM trigger does not change timbre but slightly destroys speech quality. Thus, the WER result is not better than our proposed method’; The other triggers cause significant damage to speech quality, resulting in lower TCR and WER. Generally, the methods that keep timbre and content unchanged can deceive automatic deep detection and gain good stealthiness.
Table 4: The stealthiness evaluation with all triggers.
[…]
BadNets 92.8 1.46
PIBA 78.7 19.5
DABA 56.7 23.1
Ultrasonic 86.7 10.5
PBSM 93.6 2.67
VSVC 0.941 1.24
RSRT(Stretch) 98.7 1.05
RSRT(Squeeze) 97.6 1.25
Figure 5: The mel spectrogram visualization of different poison samples with mentioned triggers. The (b) and (c) show our proposed triggers. The (d)-(e) shows the audio clip backdoor attacks. The (e) and (f) shows the speech components backdoor attacks.
4.5 Ablation Study
In this section, we discuss the effect of hyper-parameters in our attack. The key hyper-parameters are the ratios of rhythm transformation and poisoning number. We can control the selected frames and copy times in the RSRT operations to produce different lengths of poisoned utterances. The RSRT ratio is equal to the stretched or squeezed length divided by the origin length. Each experiment is repeated three times to reduce the effect of randomness.
Effects of Poisoning Number: We replace the poisoning rate with the poisoning number because our proposed trigger is effective and needs a very low number of poisoning samples. The poisoning rate corresponding to 50 samples is $0.22\%$. As shown in Table 5, the ASR increases with the increase in the number of poisoning number in general. It is indicated that in cases of low poisoning rates, compared to squeezing.
Effects of Rhythm Transformation Ratio: The best effectiveness is shown in the set of $0.5$ and $1.5$ rhythm transformation ratios. With squeezing, a relatively low poisoning number of 150 can achieve an excellent attack success rate of $99.97\%$. With stretching, a relatively low poisoning number of 150 can achieve an excellent attack success rate of $99.10\%$. The squeezing ASR values under each poisoned number are obviously higher than the stretching ones. This also indicates that squeezing behaves better than stretching.
Table 5: Ablation study. The ASR with different hyper-parameters
Poisoned number
RSRT ratio 50 100 150 200
Squeeze(1/2) 87.56 93.45 99.97 1.0
Squeeze(2/3) 89.92 95.52 97.75 1.0
Stretch(1.2) 10.67 65.67 78.91 86.75
Stretch(1.5) 86.72 91.45 98.91 1.0
Stretch(2.0) 77.99 87.85 99.10 1.0
5 Conclusion
The paper proposed a speech backdoor attack method, combining mainly VAD, RSRT and neural vocoder. This method achieves very high ASR while maintaining a very low poisoning rate. The proposed trigger can avoid two kinds of main detection by speaker verification system and automatic speech recognition and gains excellent stealthiness and speech quality. The experiments demonstrate the excellent performance of efficiency and stealthiness of speech backdoor attacks with our method. We think that changing the rhythm or speech components of speech is an exploratory new approach to speech backdoor attacks.